OpenStudy (anonymous):

Write the equation of a line that is perpendicular to y=4 and goes through (-2,0) How do I solve this? If the slope is zero, wouldn't the x cancel out? How do I cancel out the y?

OpenStudy (anonymous):

m=4 a slope perpendicular to it is; m=-1/4 given (-2,0); solve using point slope form \[(y-y _{1})=m(x-x _{1})\]substitute the given ordered pairs (y-0)=-1/4(x+2) y =(-1/4)x - 1/2 if the slope is zero, therefor, 0 value of x. What will remain is value of y-intercept(b) which is -1/2 I hope i/m getting a point. . . .